Find the square of the following numbers. 
35
Find the square of the following numbers. 
35

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(35)^2
= (30+5 )^2
= (30)^2+ (5)^2+ 2×30×5 [Since, (a+b)^2= a^2+b^2+2ab]​
= 900+25+300
= 1225

Thanks